# Segment Anything in 3D Gaussians

### v2

@@ 4.3 Qualitative results
Figure 7: Collision mesh for segmented 3D objects using QuickHull algorithm.
@@
Figure 8: The influence of different view numbers on LERF-fern.
4.4 Ablation Study
Figure 9: The influence of different view directions.
@@ 4.4 Ablation Study
In this section, we study the influence of selected 2D views on 3D Gaussian segmentation. Figure 8 demonstrates the segmentation quality using different numbers of views. It is noteworthy that the leaves in the “fern” scene are very tiny and challenging. Even so, with only two sparse views, our approach can achieve decent results, and the segmentation quality quickly improves when increasing the number of views. Table 4 also draws a similar conclusion: as the number of views increases, the IoU values will also increase. It is worth noting that even with sparse views (10% percentage), we can obtain relatively decent results for both two types of scenes, which also demonstrates the robustness and effectiveness of our approach. In practical applications, using sparse views (below 10% of total views) will greatly improve efficiency, although under 100% views, our method can still be completed within one minute.
Figure 9 studies the influence of view directions on final segmentation results. We choose the truck with 360-degree views. The upper row shows the result of using roughly a single direction. In this case, the background is mistakenly incorporated into the segmentation. By contrast, with the same number of views but variant directions, the approach achieves high-quality clean segmentation results. This result reveals the importance of choosing non-monotone view directions in practice usage with 360 scenes.
Gaussian Decomposition:
Figure 10 shows the visualization results for comparison. We select two representative scenes in which the junction/contact exists large-scale Gaussian points (can be seen in the 2nd column). This issue is alleviated after utilizing the proposed Gaussian Decomposition strategy. Though this idea is simple, the improvement is obvious.
@@
In the process of our method, we set a hyper-parameter $\tau$ to control the threshold of label confidence score in Section 3.5. As shown in Table 5, too large $\tau$ values will lead to the partially missing the segmented objects, and too small $\tau$ values will remain more background objects. For scenes with some views that may cover only part of the objects to be segmented (i.e. half of the truck body in the “truck” scene), a small $\tau$ value is preferred. In the interactive segmentation process, this threshold can be set manually according to the complexity of the scenes and the quality of camera views.
@@ 5 Conclusion
In this paper, we present a novel approach for object segmentation in 3D Gaussians, which performs an interactive procedure without any training process and learnable parameters. Based on a trained 3D Gaussian representation and clicked points on first view, we propose the multiview mask generation and view-wise label assignment scheme to generalize SAM to achieve 3D segmentation. In addition, we provide the Gaussian Decomposition to mitigate the boundary roughness issue of segmented masks, and the cross-view label voting approach to assign labels from different views. Extensive segmentation experiments show that our method can effectively obtain high-quality object segmentation results in 3D Gaussians, which can be easily applied for scene editing and collision detection tasks. Overall, we hope our work can inspire more future work in the area of 3D Gaussian representation.

### v3

@@ 4.4 Ablation Study
In this section, we study the influence of selected 2D views on 3D Gaussian segmentation. Fig. 9 (b) demonstrates the segmentation quality using different numbers of views. It is noteworthy that the leaves in the “fern” scene are very tiny and challenging. Even so, with only two sparse views, our approach can achieve decent results, and the segmentation quality quickly improves when increasing the number of views. Table 4 also draws a similar conclusion: as the number of views increases, the IoU values will also increase. It is worth noting that even with sparse views (10% percentage), we can obtain relatively decent results for both two types of scenes, which also demonstrates the robustness and effectiveness of our approach. In practical applications, using sparse views (below 10% of total views) will greatly improve efficiency, although under 100% views, our method can still be completed within one minute.
@@
Fig. 9 (a) study the influence of view directions on final segmentation results. We choose the truck with 360-degree views. The upper row shows the result of using roughly a single direction. In this case, the background is mistakenly incorporated into the segmentation. By contrast, with the same number of views but variant directions, the approach achieves high-quality clean segmentation results. This result reveals the importance of choosing non-monotone view directions in practice usage with 360 scenes.
@@
In the process of our method, we set a hyper-parameter $\tau$ to control the threshold of label confidence score in Sec. 3.5. As shown in Table 5, too large $\tau$ values will lead to the partially missing the segmented objects, and too small $\tau$ values will remain more background objects. For scenes with some views that may cover only part of the objects to be segmented (i.e. half of the truck body in the “truck” scene), a small $\tau$ value is preferred. In the interactive segmentation process, this threshold can be set manually according to the complexity of the scenes and the quality of camera views.
5 Discussions and Limitations
Through experiments, we found that our method does not perform well in objects where 3D Gaussians are very sparse, such as the LLFF-room [22] scene. The Gaussians of the table are notably sparse; even worse, the Gaussians representing the table surface are across different objects. Though our Gaussian Decomposition can somewhat remedy the boundary issue, it still suffers from extremely sparse points and leaves holes in 3D segmentation. We believe this limitation can be alleviated by future research in a more structured 3D-GS representation, yielding more accurate results.
Figure 9: Ablation on view selection.
@@ 6 Conclusion
We address the issue that rough/incomplete boundaries in segmenting 3D-GS and propose a novel Boundary-enhanced segmentation pipeline. Given input prompts on the first rendering view, our approach automatically generates multi-view masks and achieves consistent 3D segmentation via the proposed assignment strategy. Our Gaussian Decomposition module can effectively mitigate the boundary roughness issue of segmented objects resulting from the inherent geometry structure in 3D-GS. Extensive segmentation experiments show that our method can effectively obtain high-quality 3D object segmentation without boundary issues, and different scene-editing tasks demonstrate that our method can be easily applied to downstream applications. Overall, we hope our work can inspire more future work in the area of 3D Gaussian representation.